# Gaussians on their Way: Wasserstein-Constrained 4D Gaussian Splatting with State-Space Modeling

### v2

@@ 2.2 Dynamic Scene State Estimation
Recent advances in dynamic scene reconstruction have explored various approaches for tracking and modeling temporal changes. In object tracking, methods like SORT [53] and SLAM systems [54] have provided robust frameworks for state estimation. The integration of learning-based approaches has enhanced these methods, particularly in handling complex scenarios with limited observations [55, 56].
Recent works have explored optimal transport and probabilistic approaches for dynamic scenes. GaussianCube [57] models scenes with probabilistic distributions for robust deformation handling, while Shape of Motion [58] leverages geometric transformations for temporal coherence. KFD-NeRF [59] applies Kalman filtering to NeRF but is constrained by its discrete point representation, while optimal transport has shown promise in improving dynamic NeRF convergence [60]. Our work differs by leveraging Gaussian Splatting, which models dynamic elements as full probability distributions in the Wasserstein space. This approach captures the geometric nature of distribution transformations, which is especially beneficial for scenes with significant deformations or rapid motions.
@@ 3 Method
Our framework integrates three key components for dynamic scene rendering (Figure 1). First, a state-space updating mechanism with neural Gaussian deformation estimates motion patterns. Second, Wasserstein distance regularization ensures smooth parameter updates. Third, we model Gaussian dynamics under Wasserstein geometry for accurate motion prediction with intrinsic physical plausibility. We detail each component in the following sections.
@@ 3.1.3 Merging: Kalman-like State Updating
The Kalman Filter [61] is a recursive algorithm that estimates the state of a linear dynamical system from a series of noisy observations. It combines prior predictions with new measurements to produce optimal state estimates, accounting for uncertainties in both the process and the observations. In our context, our prediction and observation are the Gaussian distributions themselves. The counterbalancing of prior predictions and new observations allows for robust tracking of the Gaussian states over time, enabling accurate rendering of dynamic scenes. We directly apply the updated equations of the Kalman Filter to merge the predicted Gaussian $\mathcal{N}_{t}^{Ob}$ distributions with the observed data $\mathcal{N}_{t}^{P}$ to obtain the updated Gaussian distributions $\hat{\mathcal{N}}_{t}$:
@@ 4.1 Training Settings
DyNeRF [26] 29.58 - 0.080 0.015
StreamRF [27] 28.16 0.850 0.310 8.50
HyperReel [30] 30.36 0.920 0.170 2.00
NeRFPlayer [29] 30.69 - 0.110 0.05
K-Planes [31] 30.73 0.930 0.070 0.10
MixVoxels [32] 30.85 0.960 0.210 16.70
@@ 4.2 Experimental Validation and Analysis
For real-world evaluation on the Plenoptic Video Dataset, we achieve 31.62 dB PSNR and 0.940 SSIM at 37 FPS (Table 2). Figure 4 shows our method’s effectiveness in handling complex non-rigid deformations.
@@ 4.3.2 Comparison with Baseline Methods
DyNeRF 24.58 0.9240 0.0807 29.31 0.9472 0.0492 28.02 0.9395 0.0646 29.58 0.9490 0.0523
StreamRF 23.16 0.9150 0.0910 28.31 0.9372 0.0592 27.02 0.9295 0.0746 28.16 0.9350 0.0623
HyperReel 25.36 0.9354 0.0735 30.11 0.9551 0.0401 29.63 0.9433 0.0536 30.36 0.9520 0.0423
NeRFPlayer 24.69 0.9283 0.0810 29.69 0.9451 0.0510 28.69 0.9383 0.0610 30.69 0.9483 0.0510
K-Planes 24.58 0.9520 0.0824 32.50 0.9713 0.0362 28.12 0.9489 0.0662 40.05 0.9934 0.0322
4D-GS 40.02 0.9155 0.1056 37.53 0.9336 0.0580 32.71 0.8876 0.1034 40.62 0.9591 0.0600
Def-3D-Gauss 41.54 0.9873 0.0234 38.10 0.9951 0.0052 42.06 0.9867 0.0144 41.01 0.9953 0.0093
4D-Rotor-Gauss 34.25 0.9620 0.0480 39.26 0.9670 0.0380 33.33 0.9570 0.0420 34.25 0.9620 0.0480
Ours 39.06 0.9863 0.0244 40.77 0.9941 0.0048 41.52 0.9857 0.0154 42.79 0.9943 0.0103
@@
DyNeRF 28.58 0.9384 0.0607 29.58 0.9439 0.0487 29.58 0.9401 0.0585 29.58 0.9397 0.0528
StreamRF 27.16 0.9284 0.0707 28.16 0.9339 0.0587 28.16 0.9301 0.0685 28.16 0.9297 0.0628
HyperReel 29.36 0.9484 0.0507 30.36 0.9539 0.0387 30.36 0.9501 0.0485 30.36 0.9497 0.0428
NeRFPlayer 29.69 0.9383 0.0610 30.69 0.9451 0.0510 30.69 0.9483 0.0510 30.69 0.9479 0.0510
K-Planes 33.10 0.9695 0.0331 33.43 0.9737 0.0343 33.09 0.9793 0.0310 31.11 0.9708 0.0402
4D-GS 25.10 0.9384 0.0607 21.93 0.9539 0.0487 38.11 0.9301 0.0785 34.23 0.9297 0.0828
Def-3D-Gauss 33.07 0.9794 0.0183 44.16 0.9933 0.0098 44.62 0.9951 0.0063 37.72 0.9897 0.0126
4D-Rotor-Gauss 25.24 0.9570 0.0530 39.26 0.9595 0.0505 38.89 0.9620 0.0480 33.75 0.9595 0.0505
Ours 34.74 0.9784 0.0193 43.66 0.9940 0.0088 41.24 0.9945 0.0073 37.91 0.9887 0.0136
@@ 4.4.1 Effect of the State Consistency Filter
Our Filter improves motion estimation by reducing noise in observed flow (left) and producing clearer motion patterns (middle). The residual map (right) shows minimal differences between observation and prediction after training, validating that our Filter successfully balances physical consistency with observational accuracy.

### v3

@@ 2.2 Dynamic Scene State Estimation
Recent advances in dynamic scene reconstruction have explored various methods for tracking and modeling temporal changes. Traditional approaches, such as SORT [8] and SLAM systems [13], provide robust state estimation frameworks. Learning-based techniques further enhance these methods, particularly in handling complex scenarios with limited observations [54, 48].
Recent studies have investigated probabilistic and optimal transport-based approaches for dynamic scenes. GaussianCube [70] models scenes using probabilistic distributions to handle deformations robustly, while Shape of Motion [57] leverages geometric transformations for temporal coherence. KFD-NeRF [69] applies Kalman filtering to NeRF but is limited by its discrete point representation, which lacks the geometric structure of optimal transport. In contrast, our method employs 3D Gaussian splatting in Wasserstein space, integrating both position and covariance (i.e., shape) as fundamental components of each Gaussian distribution. This enables smooth updates by evolving distributions across frames on a Wasserstein manifold. Optimal transport has also shown promise in improving dynamic NeRF convergence in OTDNeRF [47], which applies unconstrained Wasserstein transport to rendered images or latent spaces, our approach directly considers the Wasserstein distance between dynamic Gaussians.
Our method distinguishes itself by leveraging Gaussian splatting to model dynamic elements as full probability distributions in Wasserstein space. This formulation captures the geometric nature of distribution transformations, making it particularly effective for scenes with significant deformations or rapid motion.
@@ 3 Method
Our framework integrates three key components for dynamic scene rendering (Figure 1). Sec. 3.1 provides a simple Euclidean-based framework for building the state space for Gaussians across frames, assuming linear motion for Gaussians. However, handling covariance as separate elements is flawed; covariance must be treated as a unified entity. Thus, in Sec. 3.2, we introduce Wasserstein distance as a proper distributional metric and the Wasserstein distance algorithm for regularization, and in Sec. 3.3, we extend it to the Wasserstein geometry on the Gaussian manifold, enabling smooth, physically intuitive Gaussian dynamics.
@@ 3.1.3 Merging: Kalman-like State Updating
The Kalman Filter [25] is a recursive algorithm that estimates the state of a linear dynamical system from a series of noisy observations. It combines prior predictions with new measurements to produce optimal state estimates, accounting for uncertainties in both the process and the observations. Learning from Kalman-like filters, we designed a fusion mechanic to merge two Gaussians: one predicted by the Wasserstein dynamic on previous stages $(t-2,t-1)$ and the other newly observed by the network at $t$. We do not adopt distinct process/measurement covariances for sensor noise of location vectors. Instead, we treat each Gaussian as a whole element and merely inherit “Kalman gain” as weights to take the balance from prediction and observation. In other words, there is no noise distribution for prediction or observation but a simple merging of the two Gaussians.
In our context, our prediction and observation are the Gaussian distributions themselves. The counterbalancing of prior predictions and new observations allows for robust tracking of the Gaussian states over time, enabling accurate rendering of dynamic scenes. We directly apply the updated equations of the Kalman Filter to merge the predicted Gaussian $\mathcal{N}_{t}^{Ob}$ distributions with the observed data $\mathcal{N}_{t}^{P}$ to obtain the updated Gaussian distributions $\hat{\mathcal{N}}_{t}$:
@@ 3.3 Modeling Gaussian Dynamics with Wasserstein Geometry
Gaussian distributions form a nonlinear manifold endowed with a natural geometry structure, making the Euclidean metric inadequate for capturing its characteristics. Moreover, it cannot plausibly capture dynamic Gaussian evolution in a physically meaningful manner. Wasserstein geometry, rooted in Optimal Transport, provides an intrinsic metric to depict the space of Gaussians. This is common sense in information geometry. However, no previous work gave explicit and differentiable algorithms of Wasserstein geometry on Gaussians and involved such powerful tools in 3D/4D Gaussian Splatting. We provide the Wasserstein Log/Exp algorithm for dynamic prediction. These advanced designs guarantee more natural, stable, coherent 4D dynamic Gaussians.
@@ 4.1 Training Settings
We establish correspondences via fixed batch indexing, where the $i$-th Gaussian in frame $t-1$ matches the $i$-th in frame $t$. We also retain 3DGS’s cloning/splitting, but delay state-consistency updates until Gaussian counts stabilize (after 3k iterations) and predicted-observed centers satisfy $\|\mu_{\text{pred}}-\mu_{\text{obs}}\|<0.1\sigma$. Gaussians with large discrepancies ($>3\sigma$) revert to standard 3DGS optimization before re-engagement, avoiding topological conflicts.
@@
DyNeRF [31] 28.31 0.9307 0.070 0.011
StreamRF [30] 27.97 0.740 0.390 7.30
HyperReel [5] 29.82 0.810 0.320 1.60
NeRFPlayer [52] 30.11 0.94 0.139 0.03
K-Planes [17] 30.73 0.930 0.141 0.10
MixVoxels [56] 30.85 0.944 0.210 16.70
@@ 4.2 Experimental Validation and Analysis
For real-world evaluation on the Plenoptic Video Dataset, we achieve 32.79 dB PSNR and 0.945 SSIM at 37 FPS (Table 2). Figure 4 shows our method’s effectiveness in handling complex non-rigid deformations.
@@ 4.3.2 Comparison with Baseline Methods
DyNeRF 26.28 0.9245 0.1030 30.54 0.9472 0.0715 28.72 0.9395 0.0869 30.28 0.9454 0.0746
StreamRF 24.93 0.8357 0.3144 29.08 0.8582 0.3026 27.79 0.8492 0.3150 28.77 0.8557 0.3057
HyperReel 26.42 0.9122 0.1914 31.17 0.9319 0.1680 30.69 0.9119 0.1715 31.29 0.9274 0.1702
NeRFPlayer 26.12 0.9289 0.1390 30.92 0.9468 0.0994 29.92 0.9383 0.1193 31.93 0.9493 0.0929
K-Planes 26.64 0.9426 0.0679 33.31 0.9519 0.0317 30.38 0.9395 0.0527 34.11 0.9740 0.0297
4D-GS 30.76 0.9516 0.0439 34.75 0.9524 0.0363 31.94 0.9697 0.0257 32.36 0.9632 0.0183
Def-3D-Gauss 32.37 0.9614 0.0345 28.93 0.9651 0.0263 32.19 0.9587 0.0255 31.84 0.9673 0.0304
4D-Rotor-Gauss 33.03 0.9651 0.0446 38.04 0.9581 0.0419 32.21 0.9579 0.0456 33.03 0.9625 0.0397
Ours 34.38 0.9663 0.0274 37.17 0.9711 0.0328 32.32 0.9717 0.0254 38.19 0.9743 0.0243
@@
DyNeRF 29.98 0.9504 0.0830 30.28 0.9419 0.0710 30.28 0.9411 0.0808 30.28 0.9397 0.0751
StreamRF 27.93 0.8488 0.3141 28.93 0.8536 0.3021 28.93 0.8498 0.3119 28.93 0.8494 0.3162
HyperReel 30.42 0.9252 0.1686 31.37 0.9107 0.1566 30.12 0.9169 0.1694 31.55 0.9265 0.1657
NeRFPlayer 30.91 0.9411 0.1160 30.80 0.9451 0.1197 32.02 0.9483 0.0990 32.91 0.9479 0.0950
K-Planes 30.16 0.9501 0.0286 31.49 0.9453 0.0398 32.15 0.9489 0.0365 30.17 0.9514 0.0357
4D-GS 31.12 0.9425 0.0259 34.23 0.9599 0.0370 35.30 0.9642 0.0334 23.97 0.9612 0.0362
Def-3D-Gauss 27.79 0.9514 0.0294 34.99 0.9569 0.0339 35.45 0.9578 0.0374 32.44 0.9617 0.0227
4D-Rotor-Gauss 29.48 0.9628 0.0570 38.04 0.9611 0.0545 37.67 0.9625 0.0459 32.53 0.9665 0.0571
Ours 30.14 0.9684 0.0241 38.06 0.9712 0.0268 32.03 0.9645 0.0213 33.31 0.9787 0.0266
@@ 4.4.1 Effect of the State Consistency Filter
As shown in Figure 5, our Filter improves motion estimation by reducing noise in the observed flow (left) and producing clearer motion patterns (middle). The residual map (right) indicates minimal differences between observation and prediction after training, validating that our Filter successfully balances physical consistency with observational accuracy. The results in Table 5 confirm this improvement, with the State Consistency Filter achieving a lower AEPE compared to the baseline.